OpenStudy (anonymous):

show me how to simplify 6+4x=1/3 (6x+9)

OpenStudy (anonymous):

6+4x=1/3 (6x+9) multiply both sides by 3 3(6+4x)=(6x+9) 18 + 12x =6x+9 subtract 18 to each side 12x =6x - 9 subtract 6x from both sides 6x = -9 divide by 6 on both sides x = -9/6 = -3/2

OpenStudy (angela210793):

6+4x=2x+3

OpenStudy (angela210793):

she said just simplify actually..not solve for X.....:)

OpenStudy (anonymous):

solve for x

OpenStudy (angela210793):

Gianfranco has solved it then :)

OpenStudy (anonymous):

Ange do not get to literal about questions. If you see an equals sign it is an equation and the implication is that it needs to be solved.

OpenStudy (angela210793):

i was abt writing tht..but he/she said simplify...and i quit :P:P:P:P thnx anyway :):):)